OpenStudy (anonymous):

125(1/5)^x_4=5 how do i solve to get x=?

OpenStudy (anonymous):

\[\large 125(\frac{1}{5})^{x_4}=5\] divide by 125 u have \[\large (\frac{1}{5})^{x_4}=\frac{5}{125}=\frac{5}{5*25}=\frac{1}{25}\] does that help?

OpenStudy (anonymous):

not really sorry still confused :/ because now i have \[(1\5)^x_4=1/25\]

OpenStudy (anonymous):

should i make the 1/5 into (5/25)^x_4

OpenStudy (anonymous):

just compare the LHS and RHS in the equation i gave u

OpenStudy (anonymous):

what does \(x_4\)mean?

OpenStudy (anonymous):

is it \(\frac{x}{4}\) or \(4x\) or something else?

OpenStudy (anonymous):

x sub 4 power

OpenStudy (anonymous):

so it is just a variable then call it \(x\)

OpenStudy (anonymous):

btw not to be critical but there is not such thing as "x sub 4 power" there is \(x^4\) read "x to the 4th power" or \(x_4\) read "x four"

OpenStudy (anonymous):

i guess its x4 then :o

OpenStudy (anonymous):

if you have \[\left(\frac{1}{5}\right)^x=\frac{1}{25}\] then you can solve for \(x\) by asking what power of \(\frac{1}{5}\) would get \(\frac{1}{25}\)

OpenStudy (anonymous):

how do i get rid of the subscrpt then ? \[125 (1/5)^(x_4) = 5 \]

OpenStudy (anonymous):

\[(1/5)^(x_4)=1/25\]